### v2

# On Strong Lefschetz Property of 0-dimensional complete intersections

Abstract.
We prove that a homogeneous 0-dimensional complete intersection satisfies the Strong Lefschetz Property (SLP) in degree 1 if and only if its associated form has nonzero Hessian.
Key words and phrases:
0-dimensional complete intersections, Lefschetz properties
2020 Mathematics Subject Classification:
Primary 14M10; Secondary 13A02, 13E10
1. Statements of main results
In [4], the authors establish that a homogeneous 0-dimensional complete intersection satisfies the Strong Lefschetz Property (SLP) in degree 1 provided that its associated form has a nonzero discriminant. Here, we show that this condition can be weakened: the necessary and sufficient condition is that the Hessian of the associated form is nonzero.
Following the notation of [4], let $S=\mathbb{C}[x_{1},\ldots,x_{n}]$ be the polynomial ring with the usual grading and $n\geq 3$, and $f_{j}\in S_{d_{j}}$ be a homogeneous polynomial of degree $d_{j}\geq 2$ for $j=1,\ldots,n$ such that the ideal
$J({\bf f})=(f_{1},\ldots,f_{n})$
is a 0-dimensional complete intersection. The quotient $M({\bf f})=S/J({\bf f})$ is a graded Artinian Gorenstein algebra with socle degree
(1) $T=\sum_{j=1}^{n}d_{j}-n.$
To $\bf f$}, one can associate a homogeneous polynomial $\mathrm{A}_{\bf f}$, called associated form of $\bf f$}, in the polynomial ring $R:=\mathbb{C}[y_{1},\ldots,y_{n}]$, where $y_{i}$ are dual variables to $x_{i}$ [1, 2, 5]. More precisely, denote by $\omega\colon\mathrm{Soc}(M({\bf f}))\to\mathbb{C}$ the linear isomorphism given by the condition $\omega(\overline{\mathrm{Jac}({\bf f})})=1$, where $\mathrm{Jac}({\bf f})$ is the determinant of the Jacobian matrix of the induced polynomial mapping ${\bf f}:\mathbb{C}^{n}\to\mathbb{C}^{n}$. The associated form $\mathrm{A}_{\bf f}$ is defined by the formula
(2) $\mathrm{A}_{\bf f}(y_{1},\ldots,y_{n})=\omega((y_{1}\overline{x_{1}}+\ldots+y_{n}\overline{x_{n}})^{T}),$
where $\overline{x_{i}}\in M({\bf f})$ is the image of $x_{i}$.
For a homogeneous polynomial $F\in R$ of degree $T$, its apolar ideal of $F$ is defined by
$I_{F}=\mathrm{Ann}(F)=\{g\in S\ |\ g\circ F=0\}\subseteq S,$
where the action of $S$ on $R$ is obtained by identifying $x_{i}$ with $\partial/\partial y_{i}$. Moreover, for any finitely generated graded Artinian Gorenstein $\mathbb{C}$-algebra $S/J$ with socle degree $T$, there exists a homogeneous polynomial $F\in R_{T}$, unique up to scaling, such that $J=I_{F}$; see [6, Lemma 2.12]. Any such $F$ is called a Macaulay inverse system for $S/J$. It is known that the associated form $A_{\bf f}(y_{1},\ldots,y_{n})\in R_{T}$ is a Macaulay inverse system for the Artinian algebra $M({\bf f})$, see [1, Proposition 3.2].
In this note, we provide a necessary and sufficient condition on the ideal $J({\bf f})$ for the Artinian algebra $M({\bf f})$ to satisfy SLP in degree 1, expressed in terms of the Hessian of its associated form.
Theorem 1.
Let $M({\bf f})$ be a 0-dimensional homogeneous complete intersection. Then the strong Lefschetz property holds in degree 1 for $M({\bf f})$; that is, for a generic linear form $\ell\in S_{1}$, the multiplication map
$\ell^{T-2}:M({\bf f})_{1}\to M({\bf f})_{T-1}$
is an isomorphism, if and only if its associated form $\mathrm{A}_{\bf f}$ has nonzero Hessian.
As in [4, Remark 1.2], let $f\in S_{d}$ be a reduced homogeneous polynomial of degree $d\geq 3$. By Euler’s formula, the singular subscheme of the reduced hypersurface $X=V(f)\subseteq\mathbb{P}^{n-1}$ is defined by the gradient ideal
$J(f):=\left(f_{1}=\frac{\partial f}{\partial x_{1}},\dots,f_{n}=\frac{\partial f}{\partial x_{n}}\right).$
The graded $\mathbb{C}$-algebra $M(f):=S/J(f)$ is called the Milnor algebra of $f$. If $f$ defines a smooth hypersurface in $\mathbb{P}^{n-1}$, then $J(f)$ is generated by a regular sequence; thus, $M(f)$ is an Artinian Gorenstein graded $\mathbb{C}$-algebra with socle degree T=n(d2)T=n(d-2). We denote by
$A_{f}(y_{1},\ldots,y_{n})\in R_{T}$
the Macaulay inverse system for the Milnor algebra $M(f)$. This setting is a special case of 0-dimensional homogeneous complete intersections described above. Hence Theorem 1 immediately yields the following.
Corollary 1.
Let $f$ be a polynomial of degree $d\geq 3$ defining a smooth projective hypersurface. Then the strong Lefschetz property holds in degree 1 for the Milnor algebra $M(f)$ if and only if the associated form $A_{f}$ has nonzero Hessian.
We would like to thank A. Dimca for useful remarks. The author is supported by NSFC-12301100.
2. Proof
Consider the Veronese morphism
$v:\mathbb{P}(S_{1})\to\mathbb{P}(S_{T-1}).$
Since the Veronese variety
${\mathcal{V}}=v(\mathbb{P}(S_{1}))$
spans the whole projective space $\mathbb{P}(S_{T-1})$ and $\mathbb{P}(J({\bf f})_{T-1})$ is a proper linear subspace of $\mathbb{P}(S_{T-1})$, the set
${\mathcal{A}}=v^{-1}\left({\mathcal{V}}\cap\mathbb{P}(J({\bf f})_{T-1})\right)$
is a proper algebraic subset of $\mathbb{P}(S_{1})$.
Let
$\pi:\mathbb{P}(S_{T-1})\setminus\mathbb{P}(J({\bf f})_{T-1})\to\mathbb{P}^{n-1}.$
be the linear projection with center $\mathbb{P}(J({\bf f})_{T-1})$; it is a regular morphism. Now consider the composition
$\phi=\pi\circ v:\mathbb{P}(S_{1})\setminus{\mathcal{A}}\to\mathbb{P}^{n-1}$
which is also regular.
As indicated in the proof of Theorem 1.5 of [4], the algebra $M({\bf f})$ has SLP in degree 1 exactly when $\phi$ is an immersion at some point $\ell\in\mathbb{P}(S_{1})\setminus{\mathcal{A}}$. Indeed, take any linear form $\ell\in\mathbb{P}(S_{1})\setminus{\mathcal{A}}$. The projective tangent space to the Veronese variety $\mathcal{V}$} at the point $p=\ell^{T-1}$ is given by
(3) $T_{p}{\mathcal{V}}=\{\ell^{T-2}\ell_{1}\ |\ \ell_{1}\in S_{1}\}\subset\mathbb{P}(S_{T-1}),$
see for instance [3, Section 1]. Consequently, the kernel of the differential $d\phi_{\ell}$ consists of those $\ell_{1}\in S_{1}\setminus\mathbb{C}\ell$ for which $\ell^{T-2}\ell_{1}$ belongs to the span of $\ell^{T-1}$ and $J({\bf f})_{T-1}$.
If SLP in degree 1 fails for $M({\bf f})$, then for every linear form $\ell\in\mathbb{P}(S_{1})\setminus{\mathcal{A}}$, there exists a nonzero $\ell_{1}\in S_{1}$ such that $\ell^{T-2}\ell_{1}\in J({\bf f})_{T-1}$. (Note that such $\ell_{1}$ cannot lie in $\mathbb{C}\ell$ because $\ell^{T-1}\notin J({\bf f})_{T-1}$). Hence $\phi$ is nowhere an immersion.
Conversely, assume that SLP holds in degree 1 but $\phi$ is not an immersion at any point. Then for a generic linear form $\ell$, we can find $\ell_{1}\in S_{1}\setminus\mathbb{C}\ell$ and a scalar $a\in\mathbb{C}$ with $(\ell_{1}-a\ell)\ell^{T-2}\in J({\bf f})_{T-1}$. Then SLP forces $\ell_{1}=a\ell$, a contradiction.
The morphism $\phi$ can be written explicitly as
$\phi=[F_{1},\ldots,F_{n}]$
where $F_{1},\ldots,F_{n}\in R_{T-1}$ form a basis of the annihilator $\mathrm{Ann}(J({\bf f})_{T-1})$, defined by
$\mathrm{Ann}(J({\bf f})_{T-1})=\{F\in R_{T-1}:\ g\circ F=0\text{ for all }g\in J({\bf f})_{T-1}\}.$
In fact, identifying $S$ with $R$ via $x_{i}\longleftrightarrow y_{i}$, the polar pairing induces a Hermitian inner product on $S_{T-1}$, given by
$\langle g,F\rangle=g\circ\overline{F},$
where $\overline{F}$ denotes the complex conjugation of $F$; that is,
$\overline{F}=\sum\overline{a_{\alpha}}x^{\alpha}\quad\text{if}\quad F=\sum a_{\alpha}x^{\alpha}.$
Choose a basis $G_{1},\ldots,G_{n}$ of $\mathrm{Ann}(J({\bf f})_{T-1})$ satisfying the orthogonality condition $G_{i}\circ\overline{G_{j}}=\delta_{ij}$. Observe that $\overline{G_{1}},\ldots,\overline{G_{n}}$ span the orthogonal complement of $J({\bf f})_{T-1}$ with respect to this Hermitian inner product. Consequently, the morphism $\phi$ can be expressed as
$\phi(a)=[g_{1}(a),\ldots,g_{n}(a)],\ a=[a_{1},\ldots,a_{n}]$
where $g_{i}(a)$ are determined by
$(a_{1}x_{1}+\ldots+a_{n}x_{n})^{T-1}\equiv g_{1}(a)\overline{G_{1}}+\ldots+g_{n}(a)\overline{G_{n}}\mod J({\bf f})_{T-1}.$
Applying the polar pairing with $G_{i}$ to both sides and using the fact that $G_{i}\circ J({\bf f})_{T-1}=0$, we obtain
$g_{i}(a)=G_{i}\circ(a_{1}x_{1}+\ldots+a_{n}x_{n})^{T-1}=(T-1)!G_{i}(a).$
Thus it suffices to show that the basis $F_{1},\ldots,F_{n}$ can be chosen algebraically independent over $\mathbb{C}$.
Recall that $J({\bf f})=\mathrm{Ann}(\mathrm{A}_{\bf f})$. By [6, Lemma 2.15], we have $J({\bf f})_{T-1}=\mathrm{Ann}(\mathrm{A}_{\bf f})_{T}:S_{1}$, where
$\mathrm{Ann}(\mathrm{A}_{\bf f})_{T}:S_{1}=\{g\in S_{T-1}:gS_{1}\subseteq\mathrm{Ann}(\mathrm{A}_{\bf f})_{T}\}.$
Since
$(x_{i}g)\circ\mathrm{A}_{\bf f}=0\Longleftrightarrow g\circ\left(\frac{\partial\mathrm{A}_{\bf f}}{\partial y_{i}}\right)=0,\ 1\leq i\leq n,$
it follows that $\mathrm{Ann}(J({\bf f})_{T-1})$ is spanned by the first-order partial derivatives of the associated form. These derivatives are algebraically independent if and only if the Hessian of $\mathrm{A}_{\bf f}$ is nonzero. The proof is now complete.