### v3

# On Strong Lefschetz Property of 0-dimensional complete intersections

Abstract.
We prove that a homogeneous 0-dimensional complete intersection satisfies the Strong Lefschetz Property (SLP) in degree 1 if and only if its associated form has nonzero Hessian. The result is essentially known in the literature, but our proof is different compared with the previous ones.
Key words and phrases:
0-dimensional complete intersections, Lefschetz properties
2020 Mathematics Subject Classification:
Primary 14M10; Secondary 13A02, 13E10
1. Statements of main results
In [5], the authors establish that a homogeneous 0-dimensional complete intersection satisfies the Strong Lefschetz Property (SLP) in degree 1 provided that its associated form has a nonzero discriminant. Here, we show that this condition can be weakened: the necessary and sufficient condition is that the Hessian of the associated form is nonzero.
Following the notation of [5], let $S=\mathbb{C}[x_{1},\ldots,x_{n}]$ be the polynomial ring with the usual grading and $n\geq 3$, and $f_{j}\in S_{d_{j}}$ be a homogeneous polynomial of degree $d_{j}\geq 2$ for $j=1,\ldots,n$ such that the ideal
$J({\bf f})=(f_{1},\ldots,f_{n})$
is a 0-dimensional complete intersection. The quotient $M({\bf f})=S/J({\bf f})$ is a graded Artinian Gorenstein algebra with socle degree
(1) $T=\sum_{j=1}^{n}d_{j}-n.$
To $\bf f$}, one can associate a homogeneous polynomial $\mathrm{A}_{\bf f}$, called associated form of $\bf f$}, in the polynomial ring $R:=\mathbb{C}[y_{1},\ldots,y_{n}]$, where $y_{i}$ are dual variables to $x_{i}$ [1, 2, 6]. More precisely, denote by $\omega\colon\mathrm{Soc}(M({\bf f}))\to\mathbb{C}$ the linear isomorphism given by the condition $\omega(\overline{\mathrm{Jac}({\bf f})})=1$, where $\mathrm{Jac}({\bf f})$ is the determinant of the Jacobian matrix of the induced polynomial mapping ${\bf f}:\mathbb{C}^{n}\to\mathbb{C}^{n}$. The associated form $\mathrm{A}_{\bf f}$ is defined by the formula
(2) $\mathrm{A}_{\bf f}(y_{1},\ldots,y_{n})=\omega((y_{1}\overline{x_{1}}+\ldots+y_{n}\overline{x_{n}})^{T}),$
where $\overline{x_{i}}\in M({\bf f})$ is the image of $x_{i}$.
For a homogeneous polynomial $F\in R$ of degree $T$, its apolar ideal of $F$ is defined by
$I_{F}=\mathrm{Ann}(F)=\{g\in S\ |\ g\circ F=0\}\subseteq S,$
where the action of $S$ on $R$ is obtained by identifying $x_{i}$ with $\partial/\partial y_{i}$. Moreover, for any finitely generated graded Artinian Gorenstein $\mathbb{C}$-algebra $S/J$ with socle degree $T$, there exists a homogeneous polynomial $F\in R_{T}$, unique up to scaling, such that $J=I_{F}$; see [7, Lemma 2.12]. Any such $F$ is called a Macaulay inverse system for $S/J$. It is known that the associated form $A_{\bf f}(y_{1},\ldots,y_{n})\in R_{T}$ is a Macaulay inverse system for the Artinian algebra $M({\bf f})$, see [1, Proposition 3.2].
In this note, we provide a necessary and sufficient condition on the ideal $J({\bf f})$ for the Artinian algebra $M({\bf f})$ to satisfy SLP in degree 1, expressed in terms of the Hessian of its associated form.
Theorem 1.
Let $M({\bf f})$ be a 0-dimensional homogeneous complete intersection. Then the strong Lefschetz property holds in degree 1 for $M({\bf f})$; that is, for a generic linear form $\ell\in S_{1}$, the multiplication map
$\ell^{T-2}:M({\bf f})_{1}\to M({\bf f})_{T-1}$
is an isomorphism, if and only if its associated form $\mathrm{A}_{\bf f}$ has nonzero Hessian.
As in [5, Remark 1.2], let $f\in S_{d}$ be a reduced homogeneous polynomial of degree $d\geq 3$. By Euler’s formula, the singular subscheme of the reduced hypersurface $X=V(f)\subseteq\mathbb{P}^{n-1}$ is defined by the gradient ideal
$J(f):=\left(f_{1}=\frac{\partial f}{\partial x_{1}},\dots,f_{n}=\frac{\partial f}{\partial x_{n}}\right).$
The graded $\mathbb{C}$-algebra $M(f):=S/J(f)$ is called the Milnor algebra of $f$. If $f$ defines a smooth hypersurface in $\mathbb{P}^{n-1}$, then $J(f)$ is generated by a regular sequence; thus, $M(f)$ is an Artinian Gorenstein graded $\mathbb{C}$-algebra with socle degree T=n(d2)T=n(d-2). We denote by
$A_{f}(y_{1},\ldots,y_{n})\in R_{T}$
the Macaulay inverse system for the Milnor algebra $M(f)$. This setting is a special case of 0-dimensional homogeneous complete intersections described above. Hence Theorem 1 immediately yields the following.
Corollary 1.
Let $f$ be a polynomial of degree $d\geq 3$ defining a smooth projective hypersurface. Then the strong Lefschetz property holds in degree 1 for the Milnor algebra $M(f)$ if and only if the associated form $A_{f}$ has nonzero Hessian.
We were made aware that our result has appeared in prior works (see e.g. [8, Theorem 3.1],[4, Corollary 2.14 (2)]). Thus, this note gives a self-contained proof of a known result via a different approach.
2. Proof
Consider the Veronese morphism
$v:\mathbb{P}(S_{1})\to\mathbb{P}(S_{T-1}).$
Since the Veronese variety
${\mathcal{V}}=v(\mathbb{P}(S_{1}))$
spans the whole projective space $\mathbb{P}(S_{T-1})$ and $\mathbb{P}(J({\bf f})_{T-1})$ is a proper linear subspace of $\mathbb{P}(S_{T-1})$, the set
${\mathcal{A}}=v^{-1}\left({\mathcal{V}}\cap\mathbb{P}(J({\bf f})_{T-1})\right)$
is a proper algebraic subset of $\mathbb{P}(S_{1})$.
Let
$\pi:\mathbb{P}(S_{T-1})\setminus\mathbb{P}(J({\bf f})_{T-1})\to\mathbb{P}^{n-1}.$
be the linear projection with center $\mathbb{P}(J({\bf f})_{T-1})$; it is a regular morphism. Now consider the composition
$\phi=\pi\circ v:\mathbb{P}(S_{1})\setminus{\mathcal{A}}\to\mathbb{P}^{n-1}$
which is also regular.
As indicated in the proof of Theorem 1.5 of [5], the algebra $M({\bf f})$ has SLP in degree 1 exactly when $\phi$ is an immersion at some point $\ell\in\mathbb{P}(S_{1})\setminus{\mathcal{A}}$. Indeed, take any linear form $\ell\in\mathbb{P}(S_{1})\setminus{\mathcal{A}}$. The projective tangent space to the Veronese variety $\mathcal{V}$} at the point $p=\ell^{T-1}$ is given by
(3) $T_{p}{\mathcal{V}}=\{\ell^{T-2}\ell_{1}\ |\ \ell_{1}\in S_{1}\}\subset\mathbb{P}(S_{T-1}),$
see for instance [3, Section 1]. Consequently, the kernel of the differential $d\phi_{\ell}$ consists of those $\ell_{1}\in S_{1}\setminus\mathbb{C}\ell$ for which $\ell^{T-2}\ell_{1}$ belongs to the span of $\ell^{T-1}$ and $J({\bf f})_{T-1}$.
If SLP in degree 1 fails for $M({\bf f})$, then for every linear form $\ell\in\mathbb{P}(S_{1})\setminus{\mathcal{A}}$, there exists a nonzero $\ell_{1}\in S_{1}$ such that $\ell^{T-2}\ell_{1}\in J({\bf f})_{T-1}$. (Note that such $\ell_{1}$ cannot lie in $\mathbb{C}\ell$ because $\ell^{T-1}\notin J({\bf f})_{T-1}$). Hence $\phi$ is nowhere an immersion.
Conversely, assume that SLP holds in degree 1 but $\phi$ is not an immersion at any point. Then for a generic linear form $\ell$, we can find $\ell_{1}\in S_{1}\setminus\mathbb{C}\ell$ and a scalar $a\in\mathbb{C}$ with $(\ell_{1}-a\ell)\ell^{T-2}\in J({\bf f})_{T-1}$. Then SLP forces $\ell_{1}=a\ell$, a contradiction.
The morphism $\phi$ can be written explicitly as
$\phi=[F_{1},\ldots,F_{n}]$
where $F_{1},\ldots,F_{n}\in R_{T-1}$ form a basis of the annihilator $\mathrm{Ann}(J({\bf f})_{T-1})$, defined by
$\mathrm{Ann}(J({\bf f})_{T-1})=\{F\in R_{T-1}:\ g\circ F=0\text{ for all }g\in J({\bf f})_{T-1}\}.$
In fact, identifying $S$ with $R$ via $x_{i}\longleftrightarrow y_{i}$, the polar pairing induces a Hermitian inner product on $S_{T-1}$, given by
$\langle g,F\rangle=g\circ\overline{F},$
where $\overline{F}$ denotes the complex conjugation of $F$; that is,
$\overline{F}=\sum\overline{a_{\alpha}}x^{\alpha}\quad\text{if}\quad F=\sum a_{\alpha}x^{\alpha}.$
Choose a basis $G_{1},\ldots,G_{n}$ of $\mathrm{Ann}(J({\bf f})_{T-1})$ satisfying the orthogonality condition $G_{i}\circ\overline{G_{j}}=\delta_{ij}$. Observe that $\overline{G_{1}},\ldots,\overline{G_{n}}$ span the orthogonal complement of $J({\bf f})_{T-1}$ with respect to this Hermitian inner product. Consequently, the morphism $\phi$ can be expressed as
$\phi(a)=[g_{1}(a),\ldots,g_{n}(a)],\ a=[a_{1},\ldots,a_{n}]$
where $g_{i}(a)$ are determined by
$(a_{1}x_{1}+\ldots+a_{n}x_{n})^{T-1}\equiv g_{1}(a)\overline{G_{1}}+\ldots+g_{n}(a)\overline{G_{n}}\mod J({\bf f})_{T-1}.$
Applying the polar pairing with $G_{i}$ to both sides and using the fact that $G_{i}\circ J({\bf f})_{T-1}=0$, we obtain
$g_{i}(a)=G_{i}\circ(a_{1}x_{1}+\ldots+a_{n}x_{n})^{T-1}=(T-1)!G_{i}(a).$
Thus it suffices to show that the basis $F_{1},\ldots,F_{n}$ can be chosen algebraically independent over $\mathbb{C}$.
Recall that $J({\bf f})=\mathrm{Ann}(\mathrm{A}_{\bf f})$. By [7, Lemma 2.15], we have $J({\bf f})_{T-1}=\mathrm{Ann}(\mathrm{A}_{\bf f})_{T}:S_{1}$, where
$\mathrm{Ann}(\mathrm{A}_{\bf f})_{T}:S_{1}=\{g\in S_{T-1}:gS_{1}\subseteq\mathrm{Ann}(\mathrm{A}_{\bf f})_{T}\}.$
Since
$(x_{i}g)\circ\mathrm{A}_{\bf f}=0\Longleftrightarrow g\circ\left(\frac{\partial\mathrm{A}_{\bf f}}{\partial y_{i}}\right)=0,\ 1\leq i\leq n,$
it follows that $\mathrm{Ann}(J({\bf f})_{T-1})$ is spanned by the first-order partial derivatives of the associated form. These derivatives are algebraically independent if and only if the Hessian of $\mathrm{A}_{\bf f}$ is nonzero. The proof is now complete.